there are 75 questions on a college entrance examination. two points are awarded for each correct anser, and one-half point is deducted for each wrong answer. a score of 100 indicates how many correct answers, assuming that all questions are answered?
Answer by ankor@dixie-net.com(22740)

there are 75 questions on a college entrance examination. two points are awarded for each correct anser, and one-half point is deducted for each wrong answer. a score of 100 indicates how many correct answers, assuming that all questions are answered?
:
let r = no. of right answers
then
(75-r) = no. of wrong answers
:
2r -.5(75-r) = 100
2r - 37.5 + .5r = 100
2r + .5r = 100 + 37.5
2.5r = 137.5
r = 137.5/2.5
r = 55 correct answers
:
:
See if that works out 75 - 55 = 20 wrong answers
55 * 2 = 110
20*-.5 = -10
------------
total pts 100